OpenStudy (anonymous):

to the nearest dollar, the average tuition at a public four-year college was $3109 in 1997 and $3319 in 2000. find, to the nearest dollar per year, the rate at which tuition was increasing.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

3319-3109=210/3yrs=210$rate/year

OpenStudy (anonymous):

i mean $53.00 /year rate increase